Preeti has an investment that is worth 58,800 dollars and has an expected return of 11.6 percent. The investment is expected to pay her 47,100 dollars in 3 year(s) from today and X dollars in 8 year(s) from today. What is X?
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

58800=47100/1.116^3+X/1.116^8

58800=(47100*0.719461263)+(X/1.116^8)

X=(58800-33886.62549)*1.116^8

=$59944.05(Approx).